# Finding Most Shattering Minimum Vertex Cuts of Polylogarithmic Size in Near-Linear Time

### v1

@@ 10 Conclusion and Open Problems
We have presented a Monte Carlo algorithm for finding all the $k$-shredders of a $k$-vertex-connected graph in almost-linear time for fixed values of $k$ and for all $k=\mathcal{O}(\text{polylog}(n))$. This greatly improves upon the strongly-quadratic running time of the previously best known algorithm in [CT99]. In our work, we have shown how to take advantage of recent developments in local flow algorithms [FNS+20] and connectivity oracles subject to vertex failures [Kos23]. One can hope to extend our algorithms to improve the time bound of the dynamic algorithm that maintains $k$-shredders under edge insertions and deletions in [CT99]. More importantly, it would be interesting if these results could be extended to improve the vertex-connectivity augmentation algorithm suggested by [CT99] and [Jor95]. One may also hope to obtain near-linear time algorithms for finding an even-length cycle in directed graphs, as well as bypass bottlenecks in the Pfaffian orientation problem for bipartite graphs.
@@
See 2.3 Fix a balanced $k$-shredder $S$ of $G$ with partition $(C,\mathcal{R})$. Suppose we sample two edges $(x,x^{\prime})$ and $(y,y^{\prime})$ such that $x\in\mathcal{R}$ and $y$ is in a different component from $x$ in $G\setminus S$. Then, $\texttt{Shredders}(x,y)$ will list $S$. Consider Algorithm˜7.
@@
1: $G=(V,E)$ - an $n$-vertex $m$-edge undirected graph with vertex connectivity $k$
2: $\mathcal{L}$ - a list containing all balanced $k$-shredders of $G$
3: for $800k\log n$ times do
4:   independently sample two edges $(x,x^{\prime}),(y,y^{\prime})$ uniformly at random
5:   $\mathcal{L}\leftarrow\mathcal{L}\cup\texttt{Shredders}(x,y)$
6: return $\mathcal{L}$
Proof of Lemma˜2.3.
@@
Hence, with probability $1-n^{-100}$, we will eventually sample vertices $x$ and $y$ such that $x$ and $y$ are in different components of $G\setminus S$. $\texttt{Shredders}(x,y)$ will list $S$ on line 5. Furthermore, the algorithm only returns $k$-shredders because it only lists sets returned by $\texttt{Shredders}(\cdot,\cdot)$, which are guaranteed to be $k$-shredders.
Now we calculate the time complexity. We sample two edges repeatedly for a total of $800k\log n$ rounds by line 3. For each round of sampling, we call $\texttt{Shredders}(\cdot,\cdot)$ once, which runs in $\mathcal{O}(km)$ time. We conclude that the entire algorithm runs in $\mathcal{O}(k\log n)\cdot\mathcal{O}(km)=\mathcal{O}(k^{2}m\log n)$ time. ∎
@@ Appendix B Omitted Proofs from Section 4
To prove Lemma˜4.9, we first show the following slightly different lemma. This lemma gives a workable data structure needed to prune candidate $k$-shredders efficiently.
@@
where $s_{i}=\pi_{i}(S)$ is the vertex in $S$ that lies along the $i$-th path $\pi_{i}$ in $\Pi$. This $k$-tuple can be represented as a $k$-digit number in base $\ell$, inducing the lexicographical order between candidate $k$-shredders in $\mathcal{C}$. We sort the candidates by their $k$-tuple and organize the sorted list in a $|\mathcal{C}|\times k$ table (see Figure˜7) by non-decreasing order from top to bottom. This can be done via any stable sorting algorithm to sort each $k$-tuple for each of the $k$ digits, starting from the least significant digit to the most.
@@
Using this sorted table, we can detect candidate $k$-shredders that are not straddled by others. The main idea is as follows. Suppose that a candidate $k$-shredder $S$ straddles another candidate $k$-shredder $T$. Let $s_{i}=\pi_{i}(S)$ and $t_{i}=\pi_{i}(T)$. By Definition˜4.7, there exist two paths $\pi_{i},\pi_{j}\in\Pi$ (with $i<j$) such that $\delta_{\pi_{i}}(s_{i})<\delta_{\pi_{i}}(t_{i})$ and $\delta_{\pi_{j}}(s_{j})>\delta_{\pi_{j}}(t_{j})$. Without loss of generality, let $i$ be the smallest path index such that the above holds. Observe that because $i<j$ and $\delta_{\pi_{i}}(s_{i})<\delta_{\pi_{i}}(t_{i})$, the row for $S$ must be sorted above the row for $T$. On the other hand, in the $j$-th column of this table, we will see that the value in the row for $S$ ($\delta_{\pi_{j}}(s_{j})$) is greater than the value in the row for $T$ ($\delta_{\pi_{j}}(t_{j})$). This inversion is the essential marker that we use to detect straddling. To formalize, a candidate $k$-shredder on row $s$ is straddled by another candidate if and only if there exists a column of the table with top-to-bottom-ordered entries $(e_{1},\ldots,e_{|\mathcal{C}|})$ that contains an entry $e_{t}$ with $t<s$ and $e_{t}>e_{s}$, or there exists an entry $e_{t}$ with $t>s$ and $e_{t}<e_{s}$.
@@
Proof of Lemma˜4.9.
Firstly, we apply Lemma˜B.1 to eliminate candidate $k$-shredders that are straddled by other candidates in $\mathcal{O}(k|\mathcal{C}|\log|\mathcal{C}|+\ell)$ time. We are left with a set of bridges $\mathcal{B}$ and a set $\mathcal{C}$ of candidate $k$-shredders sorted by their $k$-tuple numbers in nondecreasing order. Because we have eliminated pairwise candidate straddling, the sorted list of remaining candidate $k$-shredders is well-ordered under $\preceq,\succeq$. That is, for any pair of candidate $k$-shredders $S_{i},S_{j}$ ($i<j$), we have $S_{i}\preceq S_{j}$. Hence, if $S_{j}\preceq\Gamma$ for a bridge $\Gamma$ of $\Pi$, then $S_{i}\preceq\Gamma$ for all $i<j$. The goal is to develop a procedure that takes a bridge $\Gamma$ and determines all candidate $k$-shredders straddled by $\Gamma$ in $\mathcal{O}(\mathrm{vol}(\Gamma))$ time.
@@
In [CT99], Cheriyan and Thurimella actually used radix sort to obtain the sorted table described in figure Figure˜7. Specifically, they noted that each digit in a $k$-tuple is bounded by $n$ because all simple paths from $x$ to $y$ are of length at most $n$. Hence, sorting at most $n$ $k$-digit numbers in base $n$ can be done in $\mathcal{O}(k(n+n))=\mathcal{O}(kn)$ time using radix sort. After obtaining the sorted table, they followed the remaining steps in Lemma˜4.9. This entire pruning process therefore takes $\mathcal{O}(k(n+n)+m)=\mathcal{O}(kn+m)$ time. Hence, by Lemma˜4.8, we have obtained all $k$-shredders separating $x$ and $y$. In total, the algorithm takes $\mathcal{O}(km)+\mathcal{O}(m+n)+\mathcal{O}(kn+m)=\mathcal{O}(km)$ time. ∎
@@ Appendix C Omitted Proofs from Section 5
We can assume that $|U|=k$ as otherwise $U=\{x\}$, which is caught by line 20. Firstly, we have $\delta_{\pi}(\pi(U))<\delta_{\pi}(z)$ because $U$ does not contain the endpoints of any path by line 25. The proof now becomes identical to that of Lemma˜5.9.
@@
Hence, the vertices of $P(u{\rightsquigarrow}v)$ exist as a bridge $\Gamma$ of $\Pi$ with attachment set $\{u,v,\ldots\}$. If there are no vertices in $P(u{\rightsquigarrow}v)$, then $\Gamma$ is the edge $(u,v)$. Furthermore, we have that $\delta_{\pi_{1}}(u)<\delta_{\pi_{1}}(\pi_{1}(U))$ and $\delta_{\pi_{2}}(v)>\delta_{\pi_{2}}(\pi_{2}(U))$. Therefore, $\Gamma$ straddles $U$. Lemma˜5.10 implies $U$ would have been pruned. This is checked by line 25, which implies the algorithm would have returned $\varnothing$ for $U$, giving our desired contradiction. ∎

### v2

@@ 10 Conclusion and Open Problems
We have presented a Monte Carlo algorithm for finding all the $k$-shredders of a $k$-vertex-connected graph in almost-linear time for fixed values of $k$ and for all $k=\mathcal{O}(\text{polylog}(n))$. This greatly improves upon the strongly-quadratic running time of the previously best known algorithm in [CT99]. In our work, we have shown how to take advantage of recent developments in local flow algorithms [FNS+20] and connectivity oracles subject to vertex failures [Kos23]. One can hope to extend our algorithms to improve the time bound of the dynamic algorithm that maintains $k$-shredders under edge insertions and deletions in [CT99]. More importantly, it would be interesting if these results could be extended to improve the vertex-connectivity augmentation algorithm suggested by [CT99] and [Jor95].
One may also investigate the Pfaffian orientation problem for bipartite graphs. As mentioned earlier, Hegde observed that one of the bottlenecks in the Pfaffian orientation algorithm shown by [RST99] is determining whether a special type of bipartite graph known as a brace contains a 4-shredder. However, while a brace on $n\geq 5$ vertices is guaranteed to be 3-connected by [RST99, 3.2], it is not necessarily 4-connected. It is unlikely that our algorithm can be modified to accomplish this task, as our algorithms for listing $k$-shredders heavily rely on the fact that between any two vertices, there exist $k$ openly-disjoint paths. This is no longer the case in $(k-1)$-connected graphs. Hence, whether there is a near-linear time algorithm for finding a 4-shredder in a brace remains an interesting open problem.
More generally, when $k$ is not a fixed constant, we are unaware of any polynomial time algorithm that finds a $k$-shredder in $(k-1)$-connected graphs. Note that it is possible for a $(k-1)$-connected graph to contain an exponential number of $k$-shredders. Specifically, one can construct $(k-1)$-connected graphs that admit $\Omega(2^{k-1})$ $(k-1)$-separators such that each $(k-1)$-separator can be extended to a $k$-shredder.
Figure 7: A 3-connected graph that admits $2^{3}$ 4-shredders.
For example, the 3-connected graph shown in Figure 7 admits at least $2^{3}$ many 3-separators by picking one red or blue vertex for each of the three middle edges. It can be shown by casework that each 3-separator of this form can be extended to a 4-shredder. While this construction seems to rule out the possibility of listing all $k$-shredders in general $(k-1)$-connected graphs in polynomial time, it is possible that the bipartiteness of braces can be exploited for a lower upper bound and faster algorithm. We leave the problem of removing this bottleneck to the reader.
@@
See 2.3 Fix a balanced $k$-shredder $S$ of $G$ with partition $(C,\mathcal{R})$. Suppose we sample two edges $(x,x^{\prime})$ and $(y,y^{\prime})$ such that $x\in\mathcal{R}$ and $y$ is in a different component from $x$ in $G\setminus S$. Then, $\texttt{Shredders}(x,y)$ will list $S$. Consider Algorithm 7.
@@
$G=(V,E)$- an $n$-vertex $m$-edge undirected graph with vertex connectivity $k$$\mathcal{L}$- a list containing all balanced $k$-shredders of $G$$800k\log n$times
1:independently sample two edges $(x,x^{\prime}),(y,y^{\prime})$ uniformly at random
2:$\mathcal{L}\leftarrow\mathcal{L}\cup\texttt{Shredders}(x,y)$EndFor
3:return $\mathcal{L}$
[…]
Proof of Lemma 2.3.
@@
Hence, with probability $1-n^{-100}$, we will eventually sample vertices $x$ and $y$ such that $x$ and $y$ are in different components of $G\setminus S$. $\texttt{Shredders}(x,y)$ will list $S$ on line 2. Furthermore, the algorithm only returns $k$-shredders because it only lists sets returned by $\texttt{Shredders}(\cdot,\cdot)$, which are guaranteed to be $k$-shredders.
Now we calculate the time complexity. We sample two edges repeatedly for a total of $800k\log n$ rounds by line 7. For each round of sampling, we call $\texttt{Shredders}(\cdot,\cdot)$ once, which runs in $\mathcal{O}(km)$ time. We conclude that the entire algorithm runs in $\mathcal{O}(k\log n)\cdot\mathcal{O}(km)=\mathcal{O}(k^{2}m\log n)$ time. ∎
@@ Appendix B Omitted Proofs from Section 4
To prove Lemma 4.9, we first show the following slightly different lemma. This lemma gives a workable data structure needed to prune candidate $k$-shredders efficiently.
@@
where $s_{i}=\pi_{i}(S)$ is the vertex in $S$ that lies along the $i$-th path $\pi_{i}$ in $\Pi$. This $k$-tuple can be represented as a $k$-digit number in base $\ell$, inducing the lexicographical order between candidate $k$-shredders in $\mathcal{C}$. We sort the candidates by their $k$-tuple and organize the sorted list in a $|\mathcal{C}|\times k$ table (see Figure 8) by non-decreasing order from top to bottom. This can be done via any stable sorting algorithm to sort each $k$-tuple for each of the $k$ digits, starting from the least significant digit to the most.
@@
Using this sorted table, we can detect candidate $k$-shredders that are not straddled by others. The main idea is as follows. Suppose that a candidate $k$-shredder $S$ straddles another candidate $k$-shredder $T$. Let $s_{i}=\pi_{i}(S)$ and $t_{i}=\pi_{i}(T)$. By Definition 4.7, there exist two paths $\pi_{i},\pi_{j}\in\Pi$ (with $i<j$) such that $\delta_{\pi_{i}}(s_{i})<\delta_{\pi_{i}}(t_{i})$ and $\delta_{\pi_{j}}(s_{j})>\delta_{\pi_{j}}(t_{j})$. Without loss of generality, let $i$ be the smallest path index such that the above holds. Observe that because $i<j$ and $\delta_{\pi_{i}}(s_{i})<\delta_{\pi_{i}}(t_{i})$, the row for $S$ must be sorted above the row for $T$. On the other hand, in the $j$-th column of this table, we will see that the value in the row for $S$ ($\delta_{\pi_{j}}(s_{j})$) is greater than the value in the row for $T$ ($\delta_{\pi_{j}}(t_{j})$). This inversion is the essential marker that we use to detect straddling. To formalize, a candidate $k$-shredder on row $s$ is straddled by another candidate if and only if there exists a column of the table with top-to-bottom-ordered entries $(e_{1},\ldots,e_{|\mathcal{C}|})$ that contains an entry $e_{t}$ with $t<s$ and $e_{t}>e_{s}$, or there exists an entry $e_{t}$ with $t>s$ and $e_{t}<e_{s}$.
@@
Proof of Lemma 4.9.
Firstly, we apply Lemma B.1 to eliminate candidate $k$-shredders that are straddled by other candidates in $\mathcal{O}(k|\mathcal{C}|\log|\mathcal{C}|+\ell)$ time. We are left with a set of bridges $\mathcal{B}$ and a set $\mathcal{C}$ of candidate $k$-shredders sorted by their $k$-tuple numbers in nondecreasing order. Because we have eliminated pairwise candidate straddling, the sorted list of remaining candidate $k$-shredders is well-ordered under $\preceq$,$\succeq$ , ⪰. That is, for any pair of candidate $k$-shredders $S_{i},S_{j}$ ($i<j$), we have $S_{i}\preceq S_{j}$. Hence, if $S_{j}\preceq\Gamma$ for a bridge $\Gamma$ of $\Pi$, then $S_{i}\preceq\Gamma$ for all $i<j$. The goal is to develop a procedure that takes a bridge $\Gamma$ and determines all candidate $k$-shredders straddled by $\Gamma$ in $\mathcal{O}(\mathrm{vol}(\Gamma))$ time.
@@
In [CT99], Cheriyan and Thurimella actually used radix sort to obtain the sorted table described in figure Figure 8. Specifically, they noted that each digit in a $k$-tuple is bounded by $n$ because all simple paths from $x$ to $y$ are of length at most $n$. Hence, sorting at most $n$ $k$-digit numbers in base $n$ can be done in $\mathcal{O}(k(n+n))=\mathcal{O}(kn)$ time using radix sort. After obtaining the sorted table, they followed the remaining steps in Lemma 4.9. This entire pruning process therefore takes $\mathcal{O}(k(n+n)+m)=\mathcal{O}(kn+m)$ time. Hence, by Lemma 4.8, we have obtained all $k$-shredders separating $x$ and $y$. In total, the algorithm takes $\mathcal{O}(km)+\mathcal{O}(m+n)+\mathcal{O}(kn+m)=\mathcal{O}(km)$ time. ∎
@@ Appendix C Omitted Proofs from Section 5
We can assume that $|U|=k$ as otherwise $U=\{x\}$, which is caught by line 13. Firstly, we have $\delta_{\pi}(\pi(U))<\delta_{\pi}(z)$ because $U$ does not contain the endpoints of any path by line 17. The proof now becomes identical to that of Lemma 5.9.
@@
Hence, the vertices of $P(u{\rightsquigarrow}v)$ exist as a bridge $\Gamma$ of $\Pi$ with attachment set $\{u,v,\ldots\}$. If there are no vertices in $P(u{\rightsquigarrow}v)$, then $\Gamma$ is the edge $(u,v)$. Furthermore, we have that $\delta_{\pi_{1}}(u)<\delta_{\pi_{1}}(\pi_{1}(U))$ and $\delta_{\pi_{2}}(v)>\delta_{\pi_{2}}(\pi_{2}(U))$. Therefore, $\Gamma$ straddles $U$. Lemma 5.10 implies $U$ would have been pruned. This is checked by line 17, which implies the algorithm would have returned $\varnothing$ for $U$, giving our desired contradiction. ∎